OpenStudy (anonymous):

What is square root of 80 in simplified radical form?

OpenStudy (anonymous):

square root of 80 meaning \[\sqrt{80}\]

OpenStudy (anonymous):

A. \[\sqrt[8]{10}\] B. \[\sqrt[10]{8}\] C.\[\sqrt[5]{4}\] D. \[\sqrt[4]{5}\]

OpenStudy (anonymous):

I think D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thank you ^_^